# for Unseen Language Generalization

### v2

@@ 5 Results and Discussion
Table 1: Performance of LinguAlchemy in MASSIVE dataset for unseen languages. These models have not seen these languages during fine-tuning. For languages in *, mBERT model also never seen the languages during pre-training.
@@ 5.1 LinguAlchemy Performance
To evaluate the effectiveness of our proposed technique on unseen languages, we trained mBERT and XLM-R on the MASSIVE dataset. Specifically, we train on 25 languages and test on 27 different, unseen languages. Our results are summarized in Table 1. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Additionally, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline. AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
From Table 1, it is shown that LinguAlchemy achieves better generalization for unseen languages. We observed a significant improvement for mBERT and a modest average improvement for the stronger XLM-R model. For mBERT, LinguAlchemy can significantly increase performance in truly unseen languages of am-ET, km-KH, mn-MN, in which mBERT has never seen during the pre-training stage nor fine-tuning. These findings show that LinguAlchemy can be useful in truly zero-shot settings.
We note that LinguAlchemy tends to even out the performance across all unseen languages, leading to a massive boost in weaker languages such as cy-GB or sw-KE. However, in languages where zero-shot performance is already strong, LinguAlchemy does not seem to provide benefits and in some cases degrade performance. This is more evident in XLM-R, where LinguAlchemy flattens performance to the 80-82% range, in which many of zero-shot performance is already reached that level or more. Regardless, the potential of our method is clear, showing it to be beneficial in cases where zero-shot performance is poor.
@@ 5.2 Effect of Scaling URIEL loss
We explore consistently scaling up the URIEL loss across various scaling factors. The results can be seen in Figure 3. It is important to note that, as we use the scale-invariant optimizer AdamW, we don’t have to worry about gradients becoming too large due to extremely large losses. Generally, we observe that a scaling factor of 10x slightly outperforms other scaling factors, and the performance appears to decline with higher scale factors.
@@
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we might aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor might be more cost-effective and beneficial. Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 2.
@@
Train Group Lang. Family Languages Num. Languages
1 Indo-European
af-ZA, bn-BD, ca-ES, cy-GB, da-DK, de-DE, el-GR, en-US, es-ES, fa-IR, fr-FR, hi-IN, hy-AM, is-IS, it-IT, lv-LV, nb-NO, nl-NL, pl-PL, pt-PT, ro-RO, ru-RU, sl-SL, sq-AL, sv-SE, ur-PK
26
2 Dravidian
Train Group 1 + kn-IN, ml-IN, ta-IN, te-IN
30
3 Afro-Asiatic
Train Group 2 + am-ET, ar-SA, he-IL
33
4 Sino-Tibetan
Train Group 3 + my-MM, zh-CN, zh-TW
36
Unseen Languages
sw-KE, km-KH, vi-VN, id-ID, jv-ID, ms-MY, tl-PH, ja-JP, ka-GE, ko-KR, mn-MN, th-TH, az-AZ, tr-TR, fi-FI, hu-HU
16
Table 3: Language family distribution used in the language family generalization experiment (§5.3)
@@ 5.3 Generalization Across Language Family
We investigate LinguAlchemy across language families to further analyze the generalization capabilities of BERT and XLM-R models. This experiment offers insight into how adaptable LinguAlchemy is to a variety of linguistic features. We perform our experiment by splitting the languages in MASSIVE according to their language families and train the model on a subset of language families while testing on the rest, unseen language families. We explore on including different subset of language families, as seen in Table 3.
The "others unseen" category includes additional language families not incorporated in the training set, serving as an "unseen" testbed. As illustrated in Figure 4, LinguAlchemy demonstrates generalization towards these unseen language families. Perhaps unsurprisingly, adding more languages and, importantly, diversity to the training data improves generalization performance. Interestingly, including the Afro-Asiatic language family group—comprising "am-ET," "ar-SA," and "he-IL," all of which have distinct scripts—resulted in a notable performance enhancement from train group 2 to train group 3 experiment.
The performance of both models, combined with LinguAlchemy underscores the advantage of including a broader spectrum of languages within training groups for enhanced model generalization. However, the impact of this diversity is not uniform across all language families: While some consistently benefit from the expansion of training data, others do not, indicating that merely increasing the volume of data from the same family may not necessarily improve performance. This inconsistency indicates the potential limitations within the models’ capacity to learn and generalize the linguistic features specific to certain language families. Consequently, our observation shows that the degree of generalization varies noticeably among different families, suggesting that while some may significantly profit from these models’ capabilities, others may require more tailored strategies to gain similar performance improvement.
@@ 5.4 Seen Language Performance
One drawback of LinguAlchemy that we have noticed is the sacrifice in performance for seen languages. This result is related to our findings in unseen languages, where LinguAlchemy flattens out the performance across all languages, thus being extremely beneficial to poorly performing ones. Additionally, we note that the performance of seen languages is flattened to a similar level, which, most of the time, is worse than that of standard in-language fine-tuning. The compiled results can be seen in Table 4
Based on this finding, LinguAlchemy is suitable for enhancing the performance of extremely low-resource languages where standard cross-lingual zero-shot fine-tuning does not improve performance. In cases where training data is available, normal fine-tuning is a better choice. We are exploring why LinguAlchemy does not help with seen languages and how to boost the performance of seen languages as future work. Nevertheless, our method is still beneficial in under-resourced settings where multilingual models perform poorly.
Performance mBERT XLM-R
Full FT Ours Full FT Ours
Unseen Lgs. 45.48% 64.68% 78.97% 80.43%
Seen Lgs. 84.52% 67.45% 86.45% 81.05%
Average 64.25% 66.01% 82.56% 80.62%
Table 4: Average performance of LinguAlchemy on unseen vs seen languages
@@ 6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 27 unseen languages in a 60-class intent classification task. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss, thus ensuring a more adaptable and robust model performance.
@@
LinguAlchemy enhances performance across many unseen languages in intent classification, yet it faces limitations. Performance on seen languages is less than ideal, indicating room for improvement through methods like weight freezing. Also, better generalization appears to reduce accuracy in seen languages, pointing to a need for balanced approaches. Currently, the research is limited to intent classification, and expanding to other NLP tasks could reveal more about its versatility. Moreover, the choice of URIEL features—syntax, geography, language family—is theoretically sound, as discussed in Chapter 3, but empirical tests with different features might refine the model further. Overcoming these limitations could greatly improve the generalizability and effectiveness of multilingual NLP models.
@@ Appendix C Language Family Experiment
Tables 7 and 8 provide a comprehensive analysis of language family performance across different training groups. These tables compare the accuracy percentages of the Multilingual BERT and XLM-RoBERTa models, respectively. The results displayed in the tables elucidate the models’ capabilities in generalizing from the training data to unseen languages. A clear trend that can be observed is the improvement in performance as the training groups progress from 1 to 4, which suggests that the models benefit from exposure to a wider variety of language families during training. The ’Average’ row at the bottom of each table indicates the mean accuracy across all language families, providing an insight into the overall performance enhancement achieved by each model with incremental training diversity.

### v3

@@ 5.1 LinguAlchemy Performance
Our results as shown in Table 5 reveal that LinguAlchemy excels across all languages in the MasakhaNews dataset, including those not encountered during the pretraining of mBERT (*) and XLM-R ($\dagger$). LinguAlchemy further demonstrates notable improvements on the Semantic Relatedness dataset, showing its ability to adapt to languages with distinct typological characteristics from the training corpus. In this experiment, we opted not to compare our method against the baseline used in the Semantic Relatedness paper because LaBSE is not zero-shot; it was pretrained with sentence similarity tasks contrasting our method’s conditions. Moreover, we excluded the MAD-X experiment from the MasakhaNews evaluation because MAD-X’s parameter-efficient approach differs fundamentally from our full finetuning approach. Collectively, these insights suggest that our LinguAlchemy can generalize across varied linguistic attributes.
Additionally, we applied the same procedure to MASSIVE dataset and the results are summarized in Table 6. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Furthermore, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline. AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
Model Language Category MASSIVE MasakhaNews SemRel
Base Ours Base Ours Base Ours
mBERT Low 48.43 66.93 ($\uparrow 18.50$) 46.25 72.70 ($\uparrow 26.45$) -0.26 -0.14 ($\uparrow 0.12$)
Medium 55.96 64.42 ($\uparrow 8.47$) 42.94 64.38 ($\uparrow 21.45$) 0.16 0.21 ($\uparrow 0.05$)
High 79.66 66.81 ($\downarrow 12.85$) 77.89 73.74 ($\downarrow 4.16$) 0.02 -0.03 ($\downarrow 0.05$)
XLM-R Low 78.15 80.90 ($\uparrow 2.76$) 47.86 79.31 ($\uparrow 31.46$) 0.08 0.40 ($\uparrow 0.32$)
Medium 80.31 80.15 ($\downarrow 0.16$) 55.87 75.79 ($\uparrow 19.92$) 0.49 0.37 ($\downarrow 0.12$)
High 86.82 80.86 ($\downarrow 5.96$) 68.14 73.79 ($\uparrow 5.65$) 0.26 0.41 ($\uparrow 0.15$)
Table 2: Performance comparison of mBERT and XLM-R models across different language categories and benchmarks.
From Table 6, it is shown that LinguAlchemy achieves better generalization for unseen languages. We observed a significant improvement for mBERT and a modest average improvement for the stronger XLM-R model. For mBERT, LinguAlchemy can significantly increase performance in truly unseen languages of am-ET, km-KH, mn-MN, in which mBERT has never seen during the pre-training stage nor fine-tuning. These findings show that LinguAlchemy can be useful in truly zero-shot settings. While LinguAlchemy significantly boosts performance in weaker languages such as cy-GB or sw-KE, it can occasionally degrade results in languages with already strong zero-shot performance, particularly evident in XLM-R where it tends to flatten results to the 80-82% range.
Dataset Language UNK %
MASSIVE am-ET 6.79
km-KH 3.81
vi-VN 0.35
Other languages <0.1
SemRel amh 3.43
hau 0.60
ary 0.44
Other languages <0.4
MasakhaNews pcm 0.43
eng 0.16
Other languages <0.1
Table 3: UNK percentages in different datasets, illustrating the prevalence of unknown tokens that LinguAlchemy successfully manages.
Despite the variations in performance, the potential of LinguAlchemy is particularly clear in scenarios where zero-shot performance is inherently weak. Our hypothesis is that the model indirectly leverages familiar scripts encountered during pretraining and helps its ability to effectively handle UNK tokens. Advances in models using byte-level tokenization units theoretically reduce or eliminate OOV tokens; however, our evaluations across the MASSIVE, MasakhaNews, and SemRel2024 datasets, as shown in Table 3, confirm that UNK tokens have a minimal impact, demonstrating the robustness of LinguAlchemy in such environments. For contexts where UNK token rates are high, the solution might be orthogonal to our approach, requiring further improvement in the base models or tokenizers that could later be integrated with LinguAlchemy.
@@ 5.2 Effect of Scaling URIEL loss
We explore consistently scaling up the URIEL loss across various scaling factors. The results can be seen in Figure 3. It is important to note that, due to the use of the scale-invariant optimizer AdamW, there is no concern about gradients becoming too large as a result of the high losses. Generally, we observe that a scaling factor of 10x slightly outperforms other scaling factors, and the performance appears to decline with higher scale factors.
@@
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor is more cost-effective and beneficial. Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 4.
@@ 5.3 Generalization Across Language Family
We investigate LinguAlchemy across language families to further analyze the generalization capabilities of BERT and XLM-R models. We perform our experiment by splitting the languages in MASSIVE according to their language families and train the model on a subset of language families while testing on the rest, unseen language families. We explore on including different subset of language families, as seen in the Appendix (Table 7).
As illustrated in Figure 4, LinguAlchemy demonstrates generalization towards these unseen language families. Perhaps unsurprisingly, adding more subset of diverse languages improves generalization performance. Notably, the inclusion of the Afro-Asiatic language group—consisting of languages such as “am-ET", “ar-SA", and “he-IL", each featuring unique scripts—has significantly enhanced performance from the second to the third training group iteration. This improvement underscores LinguAlchemy’s capability to adapt to scripts not presented during the initial training or fine-tuning phases, illustrating its robustness in generalizing across different scripts.
@@
The performance of both models, combined with LinguAlchemy underscores the advantage of including a broader spectrum of languages within training groups for enhanced model generalization. However, the impact of this diversity is not uniform across all language families: While some consistently benefit from the expansion of training data, others do not, and shows that merely increasing the volume of data from the same family may not necessarily improve performance. This inconsistency highlights the potential limitations within the models’ capacity to learn and generalize the linguistic features specific to certain language families. Consequently, our observation shows that the degree of generalization varies among different language families. This suggests that while some may significantly profit from these models’ capabilities, others may require more tailored strategies to gain similar performance improvement.
[…]
Table 6: Performance of LinguAlchemy in MASSIVE dataset for unseen languages. For languages in *, mBERT has never seen the languages during pre-training.
@@ 5.4 Seen Language Performance
While LinguAlchemy consistently improves performance across unseen languages, we note some inconsistencies concerning the performance of seen languages. In MASSIVE, we observe a noticeable performance drop in seen languages, while in contrast, we still see a massive gain in MasakhaNews and the performance of SemRel seems to be unaffected. The compiled results can be seen in Table 5.1.
As MasakhaNews focuses on extremely low-resource languages, we hypothesize that despite being exposed during pretraining, the models’ performance remains low even with standard fine-tuning methods. Hence, LinguAlchemycan significantly improve performance, even by  18% in the low-resource languages. For high-resource languages, traditional fine-tuning is a better choice. We are investigating why LinguAlchemydoes not help with some languages and how to enhance the performance of some seen languages as part of our future work. Nevertheless, our method still proves beneficial in under-resourced settings where multilingual models typically perform poorly.
@@ 6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 27 unseen languages in a 60-class intent classification task. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss, thus ensuring a more adaptable and robust performance. LinguAlchemy achieves a massive performance improvement on low-resource languages in multiple downstream tasks including intent classification ($\uparrow 18.50$), news classification ($\uparrow 31.46$), and semantic relatedness ($\uparrow 0.32$).
@@
LinguAlchemy enhances performance across many unseen languages in intent classification, yet it faces limitations. Performance on seen languages is less than ideal, indicating room for improvement through methods like weight freezing. Also, better generalization appears to reduce accuracy in seen languages, pointing to a need for balanced approaches. Currently, the research is limited to intent classification, and expanding to other NLP tasks could reveal more about its versatility. Moreover, the choice of URIEL features—syntax, geography, language family—is theoretically sound, as discussed in Section 3, but empirical tests with different features might refine the model further. Overcoming these limitations could greatly improve the generalizability and effectiveness of multilingual NLP models.
@@ Appendix A Languages in Dataset
Train Group Lang. Family Languages Num. Languages
1 Indo-European af-ZA, bn-BD, ca-ES, cy-GB, da-DK, de-DE, el-GR, en-US, es-ES, fa-IR, fr-FR, hi-IN, hy-AM, is-IS, it-IT, lv-LV, nb-NO, nl-NL, pl-PL, pt-PT, ro-RO, ru-RU, sl-SL, sq-AL, sv-SE, ur-PK 26
2 Dravidian Train Group 1 + kn-IN, ml-IN, ta-IN, te-IN 30
3 Afro-Asiatic Train Group 2 + am-ET, ar-SA, he-IL 33
4 Sino-Tibetan Train Group 3 + my-MM, zh-CN, zh-TW 36
Unseen Languages sw-KE, km-KH, vi-VN, id-ID, jv-ID, ms-MY, tl-PH, ja-JP, ka-GE, ko-KR, mn-MN, th-TH, az-AZ, tr-TR, fi-FI, hu-HU 16
Table 7: Language family distribution used in the language family generalization experiment (§5.3). The "others unseen" category includes additional language families not incorporated in the training set that we use as an “unseen" testbed.
@@ Appendix C Language Family Experiment
Tables 10 and 11 provide a comprehensive analysis of language family performance across different training groups. These tables compare the accuracy percentages of the Multilingual BERT and XLM-RoBERTa models, respectively. The results displayed in the tables elucidate the models’ capabilities in generalizing from the training data to unseen languages. A clear trend that can be observed is the improvement in performance as the training groups progress from 1 to 4, which suggests that the models benefit from exposure to a wider variety of language families during training. The ’Average’ row at the bottom of each table indicates the mean accuracy across all language families, providing an insight into the overall performance enhancement achieved by each model with incremental training diversity.
@@
Appendix D Appendix: Language Identification (LID) Experiments
This section presents the results of comprehensive language identification experiments performed across a variety of popular language detection models. The evaluation is detailed in two distinct tables:
Table 12 displays the performance of traditional language identification models such as LID-Fasttext, CLD3, CLD2, langid, and LangDetect across multiple languages within the MASSIVE dataset. These results illustrate the effectiveness of each model in correctly identifying the language of given text samples.
Table 13 focuses on the accuracy of multilingual language models, specifically XLM-R and mBERT, alongside adaptations using the MAD-X framework with embeddings from FastText and CLD3. This evaluation aims to show how these advanced models perform in the task of language identification, especially in comparison to more specialized LID tools.
Language LID-Fasttext CLD3 CLD2 langid LangDetect
ar-SA 94.25 86.45 81.58 91.78 94.13
bn-BD 99.72 97.52 89.57 96.93 99.76
de-DE 97.70 88.59 89.73 92.83 82.54
el-GR 99.68 96.91 99.77 99.84 99.64
en-US 98.61 79.44 93.43 93.96 87.82
es-ES 96.20 78.24 73.14 86.87 86.55
fi-FI 97.70 92.91 92.90 92.08 96.09
fr-FR 98.35 87.53 85.23 94.77 94.80
hi-IN 98.44 88.21 97.83 87.94 93.54
hu-HU 98.54 92.24 93.89 95.34 96.71
hy-AM 99.90 98.37 99.92 99.17 0.00
id-ID 87.20 65.86 73.54 72.68 89.32
is-IS 89.93 92.64 90.88 92.97 0.00
ja-JP 99.41 96.63 99.04 99.11 96.23
jv-ID 24.75 68.10 0.00 22.04 0.00
ka-GE 99.56 98.49 99.95 99.65 0.00
ko-KR 99.50 98.47 99.03 99.96 99.36
lv-LV 90.73 90.06 95.25 94.33 97.32
my-MM 99.93 96.90 99.97 0.00 0.00
pt-PT 92.17 83.42 77.39 77.74 84.05
ru-RU 99.27 84.48 82.35 83.79 91.32
vi-VN 98.41 95.85 97.26 98.62 99.53
zh-CN 97.55 98.07 84.33 99.64 0.00
zh-TW 95.76 94.19 0.03 99.31 0.00
Average 93.89 89.57 83.17 86.31 66.20
Table 12: Per language results of language identification evaluation in MASSIVE.
Language XLMR mBERT MAD-X
MAD-X
w/ FastText
MAD-X
w/ CLD3
ar-SA 79.32 78.35 75.72 71.92 67.79
bn-BD 83.25 80.23 78.61 76.36 74.95
de-DE 85.54 83.59 81.81 79.49 76.90
el-GR 85.07 81.74 80.93 79.56 78.51
en-US 88.16 86.45 85.78 83.89 83.15
es-ES 86.18 84.97 82.58 80.97 76.43
fi-FI 85.24 82.55 82.55 79.86 77.07
fr-FR 86.48 86.11 83.69 82.35 80.03
hi-IN 84.63 82.38 80.73 78.14 72.73
hu-HU 85.68 82.65 81.57 80.13 76.40
hy-AM 84.23 81.20 80.43 78.78 77.91
id-ID 86.52 84.67 82.01 76.03 69.30
is-IS 84.16 82.21 80.40 71.49 73.57
ja-JP 85.78 84.70 83.22 82.04 81.27
jv-ID 81.20 81.57 78.58 45.70 59.68
ka-GE 79.19 75.25 73.23 70.85 70.17
ko-KR 85.51 84.30 82.99 81.14 80.56
lv-LV 84.73 82.18 82.08 74.58 74.95
my-MM 82.18 78.01 78.48 76.36 74.98
pt-PT 86.35 85.27 83.59 80.56 77.77
ru-RU 86.65 83.96 83.52 81.74 75.45
vi-VN 86.48 83.32 82.52 79.72 78.61
zh-CN 85.41 85.24 84.23 53.09 52.69
zh-TW 83.73 82.55 81.27 52.79 52.45
Average 84.65 82.64 81.27 74.90 73.47
Table 13: Per language accuracy score of multilingual language models in MASSIVE.